Solution of the differential equation xdy – ydx = 0 represents a
(A) parabola (B) circle (C) hyperbola (D) straight line

Grade:

2 Answers

vaibhav patil
3 years ago
the solution of the equation is D straight line
dy/y=dx/x 
has the equation 
lny=lnx+c which is
lny=lnx+lnC
so 
y=Cx 
staight line
 
Vikas TU
3 years ago
We have,
xdy−ydx=0
⇒ 1/ydy−1/xdx=0
⇒ logy−logx=logC [On integrating]
⇒ y/x=C⇒y=Cx
Clearly, it represents a straight line passing through the origin.